1) Ayayai Corporation issues $560,000 of 9% bonds that are due in 8 years and pay interest semi-annually. At the time of issue, the market rate for such bonds is 8%.

Calculate the bonds’ issue price by using (1) factor Tables A.2 and A.4, (2) a financial calculator, or (3) Excel function PV. (Hint: Refer to Chapter 3 for tips on calculating.) (For calculation purposes, use 5 decimal places as displayed in the factor table provided and round final answer to 0 decimal places, e.g. 5,275.)

Issue price of bond $ _______

2) Pharoah Ltd. signed an instalment note on January 1, 2020, in settlement of an account payable of $30,500 owed to Mott Ltd. Pharoah is able to borrow funds from its bank at 8%, whereas Mott can borrow at the rate of 7%. The note calls for two equal payments of blended principal and interest to be made at December 31, 2020 and 2021.

Using (1) factor Table A.4, (2) a financial calculator, or (3) Excel function PV, calculate the amount of the equal instalment payments that will be made to Mott Ltd. (Hint: Refer to Chapter 3 for tips on calculating.) (For calculation purposes, use 5 decimal places as displayed in the factor table provided and round final answer to 2 decimal places, e.g. 5,275.75.)

Amount of the equal instalment to be paid ________
Answer #1

1)

Since interest is paid semiannually

Market Rate = 8 / 2 = 4%

No of Period = 8 x 2 = 16

Price = Present Value of Interest + Present Value of Pprincipal

= 560000 x 8% x 1/2 x PVAF (4%, 16) + 560000 x PVF(4%, 16)

= 22,400 x 11.6523 + 560000 x 0.5339

= 261,011.52+ 298,984

= 559,995.52

2)

The amount of the equal instalment payments that will be made to Mott Ltd = 30,500 / PVIFA ( 8 % , 2 )

= $30500 / 1.78326

= $ 17,106